OpenStudy (anonymous):

A fireman, 54.6 m away from a burning building, directs a stream jet of water from a ground level fire hose at an angle of 38.2 degrees above the horizontal. If the speed of the stream as it leaves the hose is 38.1 m/s, at what height will the stream of water strike the building? The acceleration due to gravity is 9.8 m/s^2.

OpenStudy (anonymous):

Correct answer: 26.6714 m.

OpenStudy (anonymous):

I'm thinking the first thing to do is look at the horizontal motion We have enough information to work out the time taken to travel from hose to building Then we can use that time to calculate the vertical displacement

OpenStudy (anonymous):

I found the time (1.824 second), but how do I find the y displacement?

OpenStudy (anonymous):

I tried two equations but they both didn't work because I think I assumed that is hit the wall at the highest part of the parabola

OpenStudy (anonymous):

just use y = Vy(0)t - gt^2/2

OpenStudy (anonymous):

where did you get that formula from? Is it derived or just one of the ones you memorized? I am slving for Vyo at the moment...

OpenStudy (anonymous):

it's one of the standard equations for uniformly accelerated motion in a straight line

OpenStudy (anonymous):

\[s(t)=s(0)+ut+\frac{ 1 }{ 2 }at^2\]

OpenStudy (anonymous):

hmm, I got a wrong answer, could you try and plug it in too? I got 23.6 m/s for the velocity in the y direction...

OpenStudy (anonymous):

y component of velocity would be 38.1sin38.2 i get 23.56

OpenStudy (anonymous):

never mind I just put it in wrong...

OpenStudy (anonymous):

looks right to me

OpenStudy (anonymous):

I got it, I just read the equation wrong, so all is well! Thanks again!

OpenStudy (anonymous):

ok : )